OpenStudy (anonymous):

Assume that the recursively defined sequence converges and find its limit a1= 2; a (subscript) n+1= 48 / 2+ a(subscript)n. What does this converge to? I had 24 but I think that's wrong? Is it 0 or 24?

OpenStudy (kinggeorge):

Hold on. Let me clarify parentheses. So you have \(a_1=2\) and \[a_{n+1}=\frac{48}{2}+a_n?\]or\[a_{n+1}=\frac{48}{2+a_n}?\]

OpenStudy (anonymous):

your second option, a(n+1) = [(48)/ (2+ a(subscript)n)]

OpenStudy (kinggeorge):

If you're told to assume that it converges, this is actually rather easy. Just assume that you've reached a point where it's converged already. I.e. set \(a_{n+1}=a_n=x\) and solve for \(x\). So you just want to solve\[x=\frac{48}{2+x}\]for \(x\). You should get two solutions. Can you tell me what they are?

OpenStudy (anonymous):

6, -8?

OpenStudy (kinggeorge):

Right. So one of those values will be the value it converges to. To decide which one, simply notice that since you're starting at \(a_1=2\), which is a positive number, \(a_n\) will be positive for all \(n\in\mathbb{Z}^+\). So the limit can't possibly be \(-8\), and so it must converge to \(6\). Did this make sense?

OpenStudy (anonymous):

Hmm, It makes sense yes, are you sure its that easy?

OpenStudy (kinggeorge):

If you're given that it converges, yes. Proving it converges is not quite so easy.

OpenStudy (anonymous):

Alright, can i ask you one more brief question?

OpenStudy (kinggeorge):

Sure.

OpenStudy (anonymous):

I have to determine if the following is convergent or divergent, and i have it as convergent but im not sure..it is: sigma with infinity as top bound, k=3, and the notation itself is (3)/ (sq rt of k^2 -5)...

OpenStudy (kinggeorge):

So\[\sum_{k=3}^\infty\frac{3}{\sqrt{k^2-5}}?\]

OpenStudy (anonymous):

yes sir.

OpenStudy (kinggeorge):

Well, since \(\sqrt{k^2-5}\) ~ \(k\), this is approximately \(\sum_{n=1}^\infty 1/n\) which diverges. So I would guess that it diverges. As for proving it... I might suggest trying a comparison test.

OpenStudy (kinggeorge):

Of course. For \(k\ge3\) we have \(\sqrt{k^2-5}<k\), so\[\frac{3}{\sqrt{k^2-5}}>\frac{1}{\sqrt{k^2-5}}>\frac{1}{k}\]So\[\sum_{k=3}^\infty\frac{3}{\sqrt{k^2-5}}>\sum_{n=3}^\infty\frac{1}{n}=\infty\]

OpenStudy (anonymous):

so it does indeed diverge alright, thank you

OpenStudy (anonymous):

so anything that results in infinity like that diverges?

OpenStudy (kinggeorge):

Yes. If it sums to infinity, people usually say it diverges.

OpenStudy (anonymous):

Alright thanks a bunch.

OpenStudy (kinggeorge):

No problem.